OpenStudy (anonymous):

The face of a tent is shaped like a triangle with an area of Fifteen feet squared. The area of the triangle is represented by the formula A equals one half b times h. The height of the tent is 4 feet. How wide is the base of the tent? 7.0 feet 7.5 feet 8.0 feet 8.5 feet

OpenStudy (anonymous):

7.5

OpenStudy (anonymous):

The formula is A = 1/2 * b * h Area= 15 height = 4 15 = 1/2 * b * 4 solve for b b = 15/2 = 7.5

OpenStudy (anonymous):

Let's plug these numbers into the formula. \[4+(1/2b)=10\] What times 1/2 (or .5) is equal to 6? (10-4)

OpenStudy (anonymous):

the formula is given

OpenStudy (anonymous):

Yes, it is, but @Hannnnah must solve it by plugging in the numbers.

OpenStudy (anonymous):

Thank youuu